11.00 points 000 cash; $500,000 land: t. Lent one of the shareholders $5,000 for moving costs, recelving a signed soxo-month note from the shareholder. O Corporation
Additional S 134
4. Based only on the completed tabulation, provide the following amounts a. Total assels at the end of the year b. Total liabilities at the end of the year $870,000 280,000 | Total stockholders equity at the end of the year. 「$ 590.000 S111,500 $ 134,500 d. Cash balance at the end of the year. e. Total current assets at the end of the year. 5. Compute the current ratio for the current year. (Round your answer to 2 decimal places.)
Answer #1
a.
East Hill Home Healthcare Services was organized as a corporation. Only a corporation issues shares of capital stock to its owners in exchange for their investment, as in transaction (a)
b.
Assets = Liabilties + Stockholders equity
Cash Short term investments Notes Receivable Land Building Equipment = Short term note payable Long term note payable + Common stock APIC Retained earnigns
Beg 50000 500000 100000 50000 = 100000 100000 + 20000 80000 400000
a 90000 9000 81000
b. -9000 14000 60000 15000 = 80000 +
c 3500 -3500 = +
d -18000 18000 = +
e = +
f -5000 5000 = +
total 111500 18000 5000 514000 156500 65000 = 100000 180000 + 29000 161000 400000
c.
Total assets 870000
total liabilties 280000
total stockholders equity 590000
cash balance 111500
total current assets 134500
(115000+18000+5000)
note:
see part b for all calculations